_________________ indicate the proportion of subjects that fall into each class of a frequency table. To calculate them, we divide the class frequency by the sum of all the frequencies.

A. Class frequencies

B. Frequencies

C. Standardized frequencies

D. Relative frequencies
Answer #1

Solution:

Option A) Class frequencies is not correct , since we are dividing Class frequencies by total frequency to get proportion.

Option B) Frequencies is same as Class frequencies, so this is also an incorrect option and

Option C) there is no term like Standardized frequencies, so this is also an incorrect option.

The proportion of subjects that fall into each class of a frequency table is obtained by dividing the class frequency by the sum of all the frequencies and it is called as relative frequency.

Thus correct option is: D. Relative frequencies

Example:

Classes Class Frequency Calculations Relative Frequency
0-100 3 =3/30 0.100
100-200 5 =5/30 0.167
200-300 11 =11/30 0.367
300-400 9 =9/30 0.300
400-500 2 =2/30 0.067
Total = 30